OpenStudy (anonymous):

if y=xln^3x, then dy/dx =?

OpenStudy (anonymous):

\[xln^3x\]

zepdrix (zepdrix):

Woops no, ;careful esama :O if the power is being applied to the log, we can't use that rule. Only when the power is being applied to the contents of the log is that true.

OpenStudy (anonymous):

yes @zepdrix a'm so sorry and thank you

zepdrix (zepdrix):

Uhhh so looks like you start with `product rule` yes kevin? :)

OpenStudy (anonymous):

yes you do (x)*(dy/dx ln^3x) + (ln^3x)*(dy/dx x)

zepdrix (zepdrix):

Good good. So just having trouble differentiating the log part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when do we include log?

zepdrix (zepdrix):

It's a composition of functions. We'll have to apply the chain rule. If you write it like this,\[\Large\bf\sf \frac{d}{dx}\ln^3x\quad=\quad \frac{d}{dx}(\ln x)^3\]It's a little easier to see what the outermost function is.

zepdrix (zepdrix):

See how the outermost function is the ( )^3 we start with power rule

OpenStudy (anonymous):

3(lnx)^2 * (dy/dx lnx)?

zepdrix (zepdrix):

ya looks good so far!

zepdrix (zepdrix):

Your derivative operator is d/dx not dy/dx just something to keep in mind

OpenStudy (anonymous):

oh thank you for the clarification

OpenStudy (anonymous):

(3(lnx)^2)/x

zepdrix (zepdrix):

good :)

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):